OpenStudy (anonymous):

A 40.0 kg crate is being lowered by means of a rope. Its downward acceleration is 3.80 m/s2. What is the force exerted by the rope on the crate? A 40.0 kg crate is being lowered by means of a rope. Its downward acceleration is 3.80 m/s2. What is the force exerted by the rope on the crate? @Physics

OpenStudy (anonymous):

the force exerted by the rope on the crate is upwards (tension in rope): downward force = (40)(9.8) upward force = (40)(3.8) downward force - upward force = (40)(9.8)-(40)(3.8) = 240 N